40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 27 15:03:02 2016, sisyphus, Metaphysics, SUS, Broken wire Unknown.jpeg
    Reply  Wed Jul 27 16:56:47 2016, rana, Metaphysics, SUS, Broken wire 
    Reply  Thu Jul 28 08:42:36 2016, Steve, Update, SUS, wire clamp e12342.jpg
    Reply  Fri Jul 29 15:17:23 2016, Steve, Update, SUS, wire and clamping test 43micr242grs.jpg
       Reply  Fri Jul 29 20:12:54 2016, gautam, Update, SUS, third time - no luck 
          Reply  Mon Aug 1 14:09:20 2016, Steve, Update, SUS, third time - no luck 
             Reply  Tue Aug 9 16:19:27 2016, Steve, Update, SUS, torque driver for wire clamp 
                Reply  Wed Aug 10 15:34:24 2016, Steve, Update, SUS, 6 in-lbs torque driver for wire clamp screw 
          Reply  Wed Aug 3 00:15:39 2016, gautam, Update, SUS, ETMX suspended 
             Reply  Wed Aug 3 09:26:54 2016, Lydia, Update, SUS, ETMX suspended: photos 8x
             Reply  Thu Aug 4 03:14:39 2016, gautam, Update, SUS, ETMX suspended 
                Reply  Fri Aug 5 02:17:10 2016, gautam, Update, SUS, ETMX Ruby Wire Standoff Ready for Gluing 8x
                   Reply  Fri Aug 5 04:43:09 2016, Koji, Update, SUS, ETMX Ruby Wire Standoff Ready for Gluing 
                   Reply  Fri Aug 5 18:21:28 2016, gautam, Update, SUS, ETMX Ruby Wire Standoff GLUED 11x
                      Reply  Sun Aug 7 14:53:39 2016, ericq , Update, SUS, ETMX Standoff gluing was successful ETMX_resonances.pdfETMX_SUSspectra.zip
                         Reply  Tue Aug 9 00:44:43 2016, gautam , Update, SUS, ETMY patch-up 
                            Reply  Tue Aug 9 15:27:57 2016, gautam , Update, SUS, ETMY patch-up IMG_2996.JPGIMG_2997.JPG
                               Reply  Wed Aug 10 03:08:03 2016, gautam , Update, SUS, ETMY patch-up 
                                  Reply  Wed Aug 10 23:45:03 2016, gautam , Update, SUS, ETMY re-suspended 7x
                                     Reply  Thu Aug 11 00:20:41 2016, Koji, Update, SUS, ETMY re-suspended 
                                        Reply  Thu Aug 11 11:56:40 2016, gautam, Update, SUS, ETMY re-suspended pitchBalancingWithBias.PDF
                                           Reply  Mon Aug 15 18:28:15 2016, gautam, Update, SUS, Air-bake preparation IMG_6816.JPG
                                              Reply  Tue Aug 16 08:04:49 2016, Steve, Update, SUS, ETMY had Aluminum bridge AlBridge.jpg
                                              Reply  Tue Aug 16 21:54:27 2016, gautam, Update, SUS, Air-bake - IN PROGRESS IMG_3005.JPG
                                                 Reply  Wed Aug 17 08:47:43 2016, ericq, Update, SUS, Air-bake finished 
                                                    Reply  Wed Aug 17 23:00:57 2016, gautam, Update, SUS, Air-bake of towers IMG_3006.JPG
                                                       Reply  Thu Aug 18 08:17:16 2016, Steve, Update, SUS, wire clamping preparation 800P.jpg
                                                       Reply  Thu Aug 18 14:14:20 2016, gautam, Update, SUS, Air-bake of towers - finished 
                                                          Reply  Fri Aug 19 22:51:12 2016, gautam, Update, SUS, ETMs first-contacted 
                                                             Reply  Sun Aug 21 16:23:05 2016, gautam, Update, SUS, ETMX suspended IMG_3018.JPG
                                                                Reply  Mon Aug 22 13:06:11 2016, gautam, Update, SUS, ETMY suspended IMG_3019.JPG
                                                                   Reply  Mon Aug 22 18:04:24 2016, gautam, Update, SUS, ETMY OSEMs inserted 
                                                                      Reply  Tue Aug 23 19:35:38 2016, gautam, Update, SUS, ETMY back in IFO 10x
                                                                         Reply  Thu Aug 25 21:11:43 2016, gautam, Update, SUS, ETMY back in IFO ETMY_BounceSpectra_25Aug2016.pdf
                                                                            Reply  Fri Aug 26 11:35:44 2016, gautam, Update, SUS, ETMY UL sensor problematic  ETMY_BounceSpectra_26Aug2016.pdf41.png19.png
                                                                               Reply  Fri Aug 26 14:05:38 2016, Steve, Update, SUS, ETMY UL sensor problematic  ETMY-ULhistory.png
                                                                               Reply  Fri Aug 26 14:09:46 2016, ericq, Update, SUS, ETMY UL sensor problematic  
                                                                                  Reply  Sun Aug 28 21:40:11 2016, gautam, Update, SUS, ETMY UL sensor problematic  ETMY_UL_problematic.png
                                                                                     Reply  Mon Aug 29 09:37:05 2016, gautam, Update, SUS, ETMY UL sensor problematic  satelliteBox.png
                                                                                        Reply  Mon Aug 29 14:42:06 2016, gautam, Update, SUS, ETMY Satellite box diagnostics IMG_6840.JPGIMG_6841.JPG
                                                                            Reply  Fri Aug 26 17:48:14 2016, gautam, Update, SUS, bounce mode coupling reduction ETMY_BounceSpectra_26Aug2016_1.pdf
                                                                               Reply  Mon Aug 29 21:53:14 2016, Lydia, Update, SUS, ETMY bounce mode coupling IMG_3051.JPG42.png
                                                                                  Reply  Tue Aug 30 20:26:36 2016, gautam, Update, SUS, ITMY installed IMG_3052.JPGSRM_sensor_level_comp.pngETMY_BounceSpectra_30Aug2016.pdf
                                                                                     Reply  Wed Aug 31 18:07:43 2016, Johannes, Update, SUS, ITMY free swinging 
                                                                                        Reply  Wed Aug 31 22:09:18 2016, gautam, Update, SUS, Y arm locked to Green  
                                                                                           Reply  Thu Sep 1 08:30:24 2016, Steve, Update, SUS, ITMY_UL is sick ITMY_SENSOR_UL.pngspareSatAmp.jpg
                                                                                              Reply  Tue Sep 6 20:30:56 2016, rana, Update, SUS, ITMY_UL is sick 
                                                                                                 Reply  Tue Sep 13 11:00:59 2016, Lydia, Update, SUS, ETMY UL glitch returned 
                                                                                           Reply  Thu Sep 1 19:59:22 2016, Johannes, Update, SUS, IR mode flashes in Y arm 
                                                                                              Reply  Fri Sep 2 21:09:08 2016, gautam, Update, SUS, Y arm locked in air 03.png56.png
                                                                                                 Reply  Tue Sep 6 20:45:14 2016, gautam, Update, SUS, X arm test masses back in chamber IMG_3137.JPGIMG_3143.JPGIMG_3142.JPGIMG_3148.JPG
                                                                                                    Reply  Thu Sep 8 22:12:36 2016, gautam, Update, SUS, X arm in place, locked to green, IR flashes visible 
                                                                                                       Reply  Fri Sep 9 17:50:02 2016, gautam, Update, SUS, Heavy doors on BS-PRM, ETMY chambers 
                                                                                                          Reply  Sat Sep 10 14:26:01 2016, rana, Update, General, PRM SRM alignment Osamu.JPG
                                                                                                             Reply  Mon Sep 12 17:15:22 2016, Johannes, Update, General, PRM SRM alignment IMG_3273.JPGIMG_3275.JPG
                                                                                        Reply  Thu Sep 1 19:18:14 2016, Johannes, Update, SUS, ITMX and ETMX preemptive table leveling 8x
                                                                                           Reply  Fri Sep 2 21:15:53 2016, gautam, Update, SUS, ITMX and ETMX preemptive table leveling 
                                                                                     Reply  Thu Sep 1 15:43:16 2016, rana, Update, SUS, OSEM adjustments 
                                                                                        Reply  Mon Sep 5 19:57:24 2016, Osamu, Update, SUS, OSEM adjustments 
Message ID: 12480     Entry time: Fri Sep 9 17:50:02 2016     In reply to: 12478     Reply to this: 12481
Author: gautam 
Type: Update 
Category: SUS 
Subject: Heavy doors on BS-PRM, ETMY chambers 

[steve, teng, johannes, lydia, gautam]

  • we set about doing some final checks on the Y arm while Johannes and Lydia worked on the X arm alignment
  • locked IMC, turned on Oplev HeNes for ITMY, SRM, PRM, BS and ETMY
  • I first went into the BS-PRM chamber. Traced Oplev paths for PRM and BS, checked that the beam is approximately centered on all the steering mirrors, and traced the beam with a clean beam card to make sure there was no clipping. The beams make it out of the vacuum onto the PDs, but are not centered
  • I also checked the Y arm green - the beam isn't quite centered on the periscope mirrors but I guess this has always been the case and I didn't venture to make any changes
  • Checked new PRM foil hats were secure
  • Checked the main IR beam out of the IMC, and also the IPANG beam - Steve suggested we keep track of the way this moves during pumpdown. However, I didn't quite think this through and we put the heavy door on the BS-PRM chamber before checking where the IPANG beam was on ETMY table (we later found that the beam was a tad too high. Anyways, this isn't critical, wouldve been nice to have this reference though
  • Checked that there were no tools lying around inside the chamber, and proceeded to put the heavy door on
  • Moved to ETMY table, and did much of the same as above - Oplev beam makes it successfully out off the ETM, OSEM cables aren't a risk to clipping the green input beam
  • Proceeded to put the heavy door on ETMY chamber
  • I would have liked to put the heavy door on the ITMY chamber today evening too, but while freeing the SRM from its EQ stops, I noticed that the LL and LR OSEM PD readouts are approximately 60 and 75 % of their saturation values. I think this warrants fixing (I also checked against the frame files from our last DRFPMI lock in march and the PD signals are significantly different) so we should do this before putting the heavy door on. It would also be a good idea to check the table leveling
  • The Oplev beams for ITMY and SRM make it cleanly out of the chamber so all looks good on that front
  • IR and green beams are well clear of any OSEM cables

Depending on how the X arm situation is, we will finish putting back all the heavy doors on Monday and start the pumpdown

GV Edit 11.30pm: 

  • We succeeded in locking the X arm as well, although the transmission peaked at 0.1 (but this is the high gain PD and not the QPD, and also, unlike the Y arm, the 50-50 BS splitting the transmitted light between the QPD and the high gain PD is still in place, so can't really compare with the Y arm value of 0.6)
  • To get the lock going, we had to change a bunch of things like the POX DC offset, demod phase, sign of the gain etc. It is unclear whether we are locking on the TEM00 mode, but we judged it is sufficient to close doors and pump down
  • Johannes and I centered the ETMX and ITMX OL spots on their respective QPDs. Earlier today, Johannes and Lydia had checked ITMX and ETMX OL paths, everything looks decent
  • JE piggyback edit : We also tied the upper ITMX OSEM cables to the suspension cage side using copper wire since particularly UR looked like it could slip and possibly fall down into the beam path
  • JE piggyback edit: While leveling the ITMX table, Gautam and I found that some of the screws that secure the weights were not vented. None of these were put in during this vent. We replaced them all with vented screws.
  • Rana also checked PRM and SRM alignment, all looks okay on that front - the OSEM problem I had alluded to earlier isn't really a problem, once the SRM is aligned, all the OSEMs are reasonably close to 50% of their saturation value.

Looks like on Monday, we will look to put the heavy doors on ITMY, ITMX and ETMX chambers, and begin the pumpdown

ELOG V3.1.3-